Problems with connectivity to main datacentre

Date: 
Mon, 28/10/2013 - 13:20 - 13:30

We experienced a Network outage at our main Data centre this afternoon around 13:20pm (GMT). The issue occurred on the backbone network of the Data centre and was fixed within 10 minutes of going down. This has affected numerous other customers using this data centre and is not an exclusive Network EQ issue. We shall be receiving an RCA (Root Cause Analysis) shortly from the data centre provider and will be reflecting the RCA here as soon as we receive it.

If you are still experiencing connection issues, please let us know asap.

Apologies for any inconvenience caused by this outage.

Network EQ Support Team

UPDATE 13:58PM: A Summary of the RCA we have received from our data centre provider: "Network Engineers investigated the reported issue and found a DDoS attack via the primary 10GBe LINX connection. The congestion as a result of this DDOS attack will have resulted in packet loss and additional latency to LINX destinations (primarily access networks in the UK and Europe) on this path whilst we were mitigating the attack. We have at this time, mitigated this attack are continuing to monitor the situation."